Circular No. 8186

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hn IN NGC 1448 R. Evans, Hazelbrook, N.S.W., reports his visual discovery with a 0.31-

m reflector on Aug. 25.70 UT of an apparent supernova (mag 14.1) located 30′′ east and 50′′ north of the nucleus of NGC 1448, adding that nothing was visible at this location to m_{v} ∼ 15.5 on Aug. 5 or on various CCD/photographic images of the galaxy. G. Bock, Bannockburn, Qld., reports that a CCD image taken on Aug. 25.74 shows SN 2003hn at mag ≈ 14.2 and located ≈ 47′′ east and 53′′ north of the galaxy nucleus. K. Krisciunas, Las Campanas Observatory and Cerro Tololo Interamerican Observatory (CTIO), reports that images taken by J. Espinoza (CTIO 1.3m telescope) yield a semi-accurate position for SN 2003hn: α = 3^{h}44^{m}36^{s}.1, δ = −44^{o}37′49′′ (equinox 2000.0; uncertainty 1′′-2′′); no star was visible at this location on a CTIO image from Jan. 3 (limiting mag V ≈ 19). A star (mag V ≈ 18.2) lies ∼ 10′′ east-northeast of SN 2003hn. Preliminary CTIO photometry (± 0.04 mag) on Aug. 26.32: V = 14.36, B − V = +0.21, U − B = −0.87 (aperture size 4′′; some galaxy light included, with SN 2003hn superimposed on the eastern edge of a possible H II region).

SUPERNOVA 2003ho IN ESO 235-G58 Further to IAUC 8161, L. A. G. Monard reports his discovery of an

apparent supernova (mag ≈ 18.6 ± 0.3) on a CCD image taken on July 31.796 UT, with a confirming image taken on Aug. 25.795 (and the new star at mag ≈ 18.8). The new object is located at α = 21^{h}06^{m}30^{s}.56, δ = −48^{o}07′29′′.9 (equinox 2000.0), which is 33′′ east and 15′′ south of the nucleus of ESO 235-G58. SN 2003ho is not visible on the Digitized Sky Survey (limiting red mag 20.5) or on an image taken by Monard on 2003 July 10.08 (limiting red mag 18.5).

V2573 OPHIUCHI R. W. Russell, D. K. Lynch, and D. L. Kim, Aerospace Corporation;

H. B. Hammel, Space Science Institute; and R. B. Perry, Langley Research Center, NASA, report on 3- to 13-μm spectroscopy of V2573 Oph on Aug. 21.25 UT with the the NASA Infrared Telescope Facility (+ BASS; 3′′.4diameter beam). The spectrum decreased monotonically toward longer wavelengths and was well fit by a 1000-K gray-body function. No emission features or infrared excess were evident. The magnitudes were L = 4.1 ± 0.05, M = 3.4 ± 0.05, and [N(10.2 μm)] = 2.3 ± 0.05.

2003 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green